OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 2x + y = 5 x + 3y = 0

OpenStudy (jmark):

Given 2x + y = 5....(1) x + 3y = 0 ….......(2) multiply by 3 in equation (1) an subtract from (2) =>6x+3y=15......(1) =>x+3y=0.......(2) - - - ------------------------ 5x=-15 =>x= -3 substitute x value in equation (1) =>2(-3)+y=5 =>-6+y=5 =>y=5+6 =>y=11 Learn more about Surface Area of a Cube http://goo.gl/fI0qFa